OpenStudy (anonymous):

Hi, Could someone help me with the quadratic approximation of x ln x? I've checked the solutions but still don't really understand. it says to substitute 1 + h, but once I do, I'm not certain how to differentiate that, nor the other steps. It's from problem set 3 (2A-12E). Thanks!

OpenStudy (phi):

The quadratic approximation is \[ f(x) = f(x_0) + f'(x_0)(x - x_0) + \frac{1}{2} f''(x_0) (x - x_0)^2 \] For your problem, we want to approximate the function \[ f(x) = x \ln(x) \] in the vicinity of x=1. That means we want to use \(x_0 = 1 \) in the approximation formula (which is the first 3 terms in a Taylor Series expansion). From the formula, you see we need \(x_0 \) (good, we have it, as 1), we need \( f(x_0) = f(1) \), which we can do. and we need the first and second derivatives of f(x), evaluated at x=1

OpenStudy (phi):

we can work through the formula and get the result. (we would need to use the product rule to find the derivative of x lnx ) In the posted answer, they took a slightly different approach (apparently to avoid taking the derivative of a product ) they said, let x be 1+ h then \[ f(x) = x \ln(x) \rightarrow f(x)= (1+h) \ln(1+h) \] now they used the quadratic approximation formula to approximate *only* the ln part in other words, evaluate about \( x_0=1\) we get \[ \ln(x) \approx \ln(1) + \frac{1}{x}\bigg|_{x=1} (x-1) + -\frac{1}{2}\frac{1}{x^2}\bigg|_{x=1}(x-1)^2 \\ \ln(x) \approx (x-1) - \frac{(x-1)^2}{2} \] or, if we use x= 1+h, \[ \ln(x) \approx h - h^2/2 \] using this approximation in \[ f(x) = x \ln(x) \approx (1+h)(h - h^2/2) \] multiply this out to get \[ x \ln(x) \approx h - h^2/2 +h^2 - h^3/2 \] Combine like terms, but only keep powers of 2 (h^3 is higher order than quadratic): \[ x \ln(x) \approx h +h^2/2 \] or , putting back (x-1) for h \[ x \ln(x) \approx (x-1) +\frac{(x-1)^2}{2} \]

OpenStudy (phi):

Looking at the pdf that went along with Section 25 http://ocw.mit.edu/courses/mathematics/18-01sc-single-variable-calculus-fall-2010/unit-2-applications-of-differentiation/part-a-approximation-and-curve-sketching/session-25-introduction-to-quadratic-appoximation/MIT18_01SCF10_Ses25a.pdf I see another reason they said let x= 1+ h in the pdf they show \[ \ln(1+x) \approx x + \frac{x^2}{2} \] pattern match that to immediately get \[ \ln(1+h) \approx h + \frac{h^2}{2} \] so just use that result in \[ x \ln x \approx (1+h) \left( h + \frac{h^2}{2} \right) \] expand, combine like terms, and drop all terms with exponents greater than 2. and then replace h with (x-1) to get their answer.